OpenStudy (anonymous):

What is the rational denominator of... 5x^4/sqrt(4x-6)

OpenStudy (tkhunny):

I do wish I knew what this question meant. Please show your work and give us a clue.

OpenStudy (anonymous):

\[\frac{ 5x^3 }{ \sqrt{4x-6} }\]

OpenStudy (anonymous):

rationalize the denominator

OpenStudy (tkhunny):

Okay, do it. What's preventing you from doing the two multiplications, one in the numerator and one in the denominator? Never post a problem with no work shown.

OpenStudy (anonymous):

What do I multiply them by? I am doing this class online and I dont get the tutorials. You dont need to be rude!!

OpenStudy (tkhunny):

I am never rude. I am honest and direct. This is often mistaken for rudeness. I have yet to see an online program that was particularly helpful. They may be out there. I just haven't seen them. You must multiply by whatever will make the denominator rational. \(\dfrac{3}{\sqrt{2}} = \dfrac{3}{\sqrt{2}}\cdot\dfrac{\sqrt{2}}{\sqrt{2}} = \dfrac{3\sqrt{2}}{2}\) See how that works?